# Architecting mechanisms of damage in topological metamaterials

### v2

@@ Appendix C Behaviour of additional topologically distinct lattices
C.1 Dominant lattice behaviour
@@
Figure 11: Dominant behaviour and failure of additional lattices. Snapshot of: (a) axial energy portion in the entire lattice; and, (b) axial stress fraction in the element that fails preceding the first failure event. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for RTK at: (c) SR = 3; and, (d) 30. The difference in angle is visualised through the colour directly to the right of each unit cell. Damage path for: (e) RTK; and, (f) RDP1. The damage process is visualised with a line traversing the domain. The line connects separated nodes, where fracture occured, within the lattice. For clarity, we do not show the micro-structure throughout the domain. The damage process for SR = 3, 10, 30 is plotted with green, black and blue lines and associated rectangular markers respectively (refer to (b)). Arrows indicate the direction of damage. Normalised maximum bending stress in cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (g, i) RDP1 (0.1,-0.1,0.1); and, (h, j) RDP3 (0.16,-0.04,0.16) at SR = 30. (k) $\theta$ spectrum for RDP1 and RDP3. (l) Damage path for RDP1 and RDP3 at SR = 30.
C.2 Dominant lattice failure
RTK dominant failure switches from bending to stretching as the SR reduces (Fig. 11b), with the damage path consistently deviating from the mid-plane (Fig. 11e). At SR = 10 and 30, where the fracture is dominated by bending, the damage paths deviate slightly above the pre-crack before propagating with small oscillations towards the right edge. At SR = 3, the failure becomes stretching-dominated, with larger oscillations in the damage path occurring as it propagates toward the right edge. In contrast, RDP1 fracture is always dominated by bending (Fig. 11b). At SR = 30, the bending stresses and damage path (Fig. 11g) align with the FM localising at the crack tip (Fig. 11i); the damage path can therefore be adjusted through their form. This is the case for RDP3, where $\theta_{\mathrm{1}}$ is initially positive (Fig. 11i), and the fracture initiates further towards the left of the domain than RDP1.
Appendix D Additional constraints
@@ D.1 Dominant SSS
D.2 Failure evolution for LP2 lattice with larger length
@@
Figure 13: Damage propagation for LP2 with n = 16, m = 50 at SR = 30.
D.3 Stress profile in lattices from (--,$+$,--) equivalence class
Here we provide the results for LUP1-3, all from the (--,$+$,--) equivalence class, to further highlight the influence of the SSS form on the stress that develops within the domain and around the crack tip. The $\theta$ spectrum for these lattices is provided within the main text. The stresses within the lattices map onto the SSS at the long-wavelength limit in all cases. For LUP1, only $\theta_{\mathrm{1}}$ can develop within the domain (Fig. 14(a)), and in Fig. 14(b) we see that the axial stress is within these zones. As SSS at $\theta_{\mathrm{1}}$ dominate the response, the bending stresses that are activated at the crack tip also follow the FM associated with $\theta_{\mathrm{1}}$ (Fig. 14(c)). In LUP2 the microstructure of the lattice is adjusted to reduce $\theta_{\mathrm{2}}$ at the long-wavelength limit whilst keeping the $\theta_{\mathrm{1}}$ spectrum similar to LUP1 (Fig. 14(d)). Compared with LUP1, additional axial stresses aligning with $\theta_{\mathrm{2}}$ appear (Fig. 14(e)). However, as the pre-crack only intersects the SSS zone associated with $\theta_{\mathrm{1}}$, the bending stresses around the crack tip for LUP2 remain similar (Fig. 14(f)). The initial damage path also matches LUP1, however as the damage evolves, the path for LUP2 (Fig. 14(d)) deviates from LUP1 (Fig. 14(a)). This may be due to the interaction with the $\theta_{\mathrm{2}}$ SSS, although this has not been investigated further here. $\theta_{\mathrm{1}}$ becomes positive at the long-wavelength limit for LUP3 (Fig. 14(g)). The axial stresses that develop within the domain also follow this form (Fig. 14(h)) and the bending stress about the crack tip shift in line with these changes (Fig. 14(i)).
Figure 14: Stress development in LUP lattices with pinned boundary conditions. Failure path, normalised axial and normalised bending stress for: (a-c) LUP1; (d-f) LUP2; and, (g-i) LUP3 respectively at SR = 30 with pinned boundary conditions. In (a), (d) and (g), we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows on either side of these zones. If the SSS can’t develop we only plot the direction of the SSS with a single arrow.
Appendix E Achiral lattices from the ($\mp$,$\pm$,$\pm$) equivalence classes
All lattices reported in the main text from the ($\mp$,$\pm$,$\pm$) equivalence classes had FM and SSS with forms that where mirror symmetric wrt. the plane of the crack. Here, we provide additional simulations for lattices with asymmetric forms for these mode, i.e. achiral unit cell geometries. In particular we choose to visualise: LP4 (0.1, -0.05, -0.15); RP4 (-0.1, 0.05, 0.15); LP5 (0.04, -0.25, -0.15); and, LP6 (0.04, -0.25, -0.05). All simulations are for SR = 30.

### v3

@@ Appendix E Behaviour of additional topologically distinct lattices
E.1 Dominant lattice behaviour
@@
E.2 Dominant lattice failure
RTK dominant failure switches from bending to stretching as the SR reduces (Fig. 12b), with the damage path consistently deviating from the mid-plane (Fig. 12e). At SR = 10 and 30, where the fracture is dominated by bending, the damage paths deviate slightly above the pre-crack before propagating with small oscillations towards the right edge. At SR = 3, the failure becomes stretching-dominated, with larger oscillations in the damage path occurring as it propagates toward the right edge. In contrast, RDP1 fracture is always dominated by bending (Fig. 12b). At SR = 30, the bending stresses and damage path (Fig. 12g) align with the FM localising at the crack tip (Fig. 12k); the damage path can therefore be adjusted through their form. This is the case for RDP3, where $\theta_{\mathrm{1}}$ is initially positive (Fig. 12k), and the fracture initiates further towards the left of the domain than RDP1.
Appendix F Additional constraints
@@
F.1 Failure evolution for LP2 lattice with larger length
Figure 14: Damage propagation for LP2 with n = 16, m = 50 at SR = 30. The simulations for this lattice is included in Movie S5.
@@ F.1 Failure evolution for LP2 lattice with larger length
Figure 15: Stress development in LUP lattices with pinned boundary conditions. Failure path, normalised axial and normalised bending stress for: (a-c) LUP1; (d-f) LUP2; and, (g-i) LUP3 respectively at SR = 30 with pinned boundary conditions. In (a), (d) and (g), we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows on either side of these zones. If the SSS can’t develop we only plot the direction of the SSS with a single arrow. The simulations for these lattices are included in Movie S5.
F.2 Stress profile in lattices from (-,++,-) equivalence class
Here we provide the results for LUP1-3, all from the (-,++,-) equivalence class, to further highlight the influence of the SSS form on the stress that develops within the domain and around the crack tip. The $\theta$ spectrum for these lattices is provided within the main text. The stresses within the lattices map onto the SSS at the long-wavelength limit in all cases. For LUP1, only $\theta_{\mathrm{1}}$ can develop within the domain (Fig. 15(a)), and in Fig. 15(b) we see that the axial stress is within these zones. As SSS at $\theta_{\mathrm{1}}$ dominate the response, the bending stresses that are activated at the crack tip also follow the FM associated with $\theta_{\mathrm{1}}$ (Fig. 15(c)). In LUP2 the microstructure of the lattice is adjusted to reduce $\theta_{\mathrm{2}}$ at the long-wavelength limit whilst keeping the $\theta_{\mathrm{1}}$ spectrum similar to LUP1 (Fig. 15(d)). Compared with LUP1, additional axial stresses aligning with $\theta_{\mathrm{2}}$ appear (Fig. 15(e)). However, as the pre-crack only intersects the SSS zone associated with $\theta_{\mathrm{1}}$, the bending stresses around the crack tip for LUP2 remain similar (Fig. 15(f)). The initial damage path also matches LUP1, however as the damage evolves, the path for LUP2 (Fig. 15(d)) deviates from LUP1 (Fig. 15(a)). This may be due to the interaction with the $\theta_{\mathrm{2}}$ SSS, although this has not been investigated further here. $\theta_{\mathrm{1}}$ becomes positive at the long-wavelength limit for LUP3 (Fig. 15(g)). The axial stresses that develop within the domain also follow this form (Fig. 15(h)) and the bending stress about the crack tip shift in line with these changes (Fig. 15(i)).
Appendix G Achiral lattices from the ($\mp$$\pm,$$\pm$) equivalence classes
All lattices reported in the main text from the ($\mp$$\pm,$$\pm$) equivalence classes had FM and SSS with forms that where mirror symmetric wrt. the plane of the crack. Here, we provide additional simulations for lattices with asymmetric forms for these mode, i.e. achiral unit cell geometries. In particular we choose to visualise: LP4 (0.1, -0.05, -0.15); RP4 (-0.1, 0.05, 0.15); LP5 (0.04, -0.25, -0.15); and, LP6 (0.04, -0.25, -0.05). All simulations are for SR = 30.